. What mass of (NH),SO, must be mixed with 35.0 kg of KNO3 to produce a fertilizer that has 15.1 mass % nitrogen?
Answer #1

Mass of (NH4)2SO4 = 7.20 kg

Explanation

mass of KNO3 = 35.0 kg

moles KNO3 = (mass KNO3) / (molar mass KNO3)

moles KNO3 = (35.0 kg) / (101.1032 g/mol)

moles KNO3 = 0.34618 kmol

moles N2 = moles KNO3

moles N2 = 0.34618 kmol

mass N2 = (moles N2) * (molar mass N2)

mass N2 = (0.34618 kmol) * (14.0 g/mol)

mass N2 = 4.846533 kg

Let mass (NH4)2SO4 = x kg

moles (NH4)2SO4 = (mass (NH4)2SO4) / (molar mass (NH4)2SO4)

moles (NH4)2SO4 = (x kg) / (132.1395 g/mol)

moles (NH4)2SO4 = x/132.1395 kmol

moles N2 = 2 * (moles (NH4)2SO4)

moles N2 = 2 * (x/132.1395 kmol)

moles N2 = 2x/132.1395 kmol

mass N2 = (moles N2) * (molar mass N2)

mass N2 = (2x/132.1395 kmol) * (14.0 g/mol)

mass N2 = 28x/132.1395 kg

Total mass N2 = 4.846533 kg + 28x/132.1395 kg

Total mass fertilizer = (mass KNO3) + (mass (NH4)2SO4)

Total mass fertilizer = (35.0 kg) + (x kg)

Total mass fertilizer = (35.0 + x) kg

mass % nitrogen = (Total mass N2 / Total mass fertilizer) * 100

15.1 = [(4.846533 kg + 28x/132.1395 kg) / (35.0 + x) kg)] * 100

0.151 = (4.846533 + 28x/132.1395) / (35.0 + x)

0.151 * (35.0 + x) = 4.846533 + 28x/132.1395

Solving for x, x = 7.20 kg

mass (NH4)2SO4 = x kg = 7.20 kg